# Design of supercurrent diode by vortex phase texture

### v3

@@ III.2 Two-vortex configuration
IV Rectification with vanishing phase gradient at the edge
Figure 6: (a) Schematic illustration for the phase vortex in the superconducting lead with two image vortices to impose the vanishing of the supercurrent at the edge. Phase vortex with the boundary condition in a superconductor at (b) $V^{\mathrm{L}}_{0}=+1$, (c) $V^{\mathrm{L}}_{0}=+2$, and (d) $V^{\mathrm{L}}_{0}=+3$. The opposite circulation of the currents at the boundary is such that the phases have a vanishing gradient perpendicular to the edge. In panels (b), (c), (d) we have a vector field at a given point $(x,y)$ whose components are given by $[-\sin(\phi(x,y),\cos(\phi(x,y)]$, where $\phi(x,y)$ is the total phase due to the vortex and the image vortices shown in Eqs. (26), (29), and (30). The phase pattern consistently shows that there is no gradient when passing through the edges along the $y$ direction. For completeness, the pattern associated with the sum of the vortex vector fields is reported in Appendix E. Vortex core is set at $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-23a,-9.5a)$ in (b)-(d).
Figure 7: (a) Schematic illustration of the Josephson junction. At the edges (green lines) we assume that the component of the supercurrent perpendicular to the boundary is vanishing. The analysis is performed for a vortex configuration having winding number $V^{\mathrm{L}}_{0}=+l$ with $l=1,2,3$. Evolution of the rectification amplitude $\eta$ with regard to the vortex core coordinates for different winding numbers: (b,c) $V^{\mathrm{L}}_{0}=1$, (f,g) $V^{\mathrm{L}}_{0}=2$, and (j,k) $V^{\mathrm{L}}_{0}=3$. Spatially resolved harmonics of the supercurrent: $I_{1}$ (red-solid line), $I_{2}$ (green-dotted), $J_{1}$ (blue-dashed), and $J_{2}$ (cyan-dashed) indicate the odd-parity first harmonic, the odd-parity second harmonic, the even-parity first harmonic, and the even-parity second harmonic amplitude, respectively. Longitudinal scan: $I_{1}$, $I_{2}$, $J_{1}$, and $J_{2}$ at a given $j^{L}_{y0}$ as a function of $j^{L}_{x0}$ for (d) $V^{\mathrm{L}}_{0}=1$, (h) $V^{\mathrm{L}}_{0}=2$, and (l) $V^{\mathrm{L}}_{0}=3$. Lateral scan: $I_{1}$, $I_{2}$, $J_{1}$, and $J_{2}$ at a given $j^{L}_{x0}$ as a function of $j^{L}_{y0}$ for (e) $V^{\mathrm{L}}_{0}=1$, (i) $V^{\mathrm{L}}_{0}=2$, and (m) $V^{\mathrm{L}}_{0}=3$. The gray dotted lines refer to the site of the maximal rectification and are a guide to indicate the values $I_{1}$, $I_{2}$, $J_{1}$, and $J_{2}$. We set $y^{\mathrm{L}}_{0}$ as (b)(d)(j)(l) $0.5a$ and (f)(h) $9.5a$, and $x^{\mathrm{L}}_{0}$ as (c)(e)(k)(m) $-23a$ and (k)(m) $-28a$. The aspect ratio is $\alpha=3/2$. The maximal rectification $\eta$ occurs for vortex core positions corresponding to a supercurrent with $I_{1}$, $I_{2}$, and $J_{1}$ components that are comparable in size. The sign change of $\eta$ is related to the vanishing of $I_{1}$ and the zeros of $J_{1}$ when the amplitude is comparable to $I_{1}$.
In the previous section, we did not take into account the effects of the boundary near the edge on the vortex phase texture, as we focused on a scenario where the edge was sufficiently distant from the vortex core, making its impact on the phase texture negligible. However, this assumption does not hold if the vortex is positioned close to the edge or if the size of the superconductor is comparable to the superconducting coherence length. In such cases, the requirement for the supercurrent to be zero perpendicular to the edge can lead to a notable alteration of the superconducting phase surrounding the vortex core. One effective way to impose this constraint on the supercurrent amplitude at the edge of the junction is through the method of vortex images. As illustrated in Fig. 6 (a), when considering a single vortex in a superconductor, we can introduce two additional image vortices outside the superconductor, positioned at a specific distance from the edge to ensure that the supercurrent amplitude becomes vanishing at the edge. It is important to note that the image vortices must have the same winding as the vortex within the superconductor because using an image antivortex would not nullify the supercurrent; rather, it would increase the supercurrent at the edge, since both circulations would be moving in the same direction.
For taking into account the boundary condition for the supercurrent amplitude at the edge, we introduce the image vortices that lead to a modification of the phase value at each site in the superconducting leads as given by
$\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}})\rightarrow\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}}),$ (25)
with
$\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}})$ =$=\begin{cases}\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}})+\varphi^{\mathrm{L}}_{a}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{a}})+(V^{\mathrm{L}}_{0}-1)\pi&j_{y}a>y^{\mathrm{L}}_{0}+(N_{y}-1)a/2\\ \varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}})+(V^{\mathrm{L}}_{0}/2-1)\pi&y^{\mathrm{L}}_{0}-(N_{y}-1)a/2\leq j_{y}a\leq y^{\mathrm{L}}_{0}+(N_{y}-1)a/2\\ \varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}})+\varphi^{\mathrm{L}}_{b}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{b}})+V^{\mathrm{L}}_{0}\pi&j_{y}a<y^{\mathrm{L}}_{0}-(N_{y}-1)a/2\end{cases},$ (26)
$\varphi^{\mathrm{L}}_{a}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{a}})$ $=V^{\mathrm{L}}_{a}\arg[(j_{x}a-x_{a})+i(j_{y}a-y_{a})],$ (27)
$\varphi^{\mathrm{L}}_{b}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{b}})$ $=V^{\mathrm{L}}_{b}\arg[(j_{x}a-x_{b})+i(j_{y}a-y_{b})],$ (28)
where $V^{\mathrm{L}}_{a}=V^{\mathrm{L}}_{b}=V^{\mathrm{L}}_{0}$ are the winding number of the image vortices while the core positions $\bm{r}^{\mathrm{L}}_{a,b}$ are defined as
$x^{\mathrm{L}}_{a}$ $=x^{\mathrm{L}}_{0},$
$y^{\mathrm{L}}_{a}$ $=(N_{y}-1)a+N_{y}a-y^{\mathrm{L}}_{0}-2a,$
$x^{\mathrm{L}}_{b}$ $=x^{\mathrm{L}}_{0},$
$y^{\mathrm{L}}_{b}$ $=-y^{\mathrm{L}}_{0}+a.$
We introduce the modulation factors to obtain the continuous phase texture in a superconductor. In Fig. 6 (b)-(d), we plot the phase texture $\bm{F}=[f_{x}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),f_{y}(\bm{j},\bm{r}^{\mathrm{L}}_{0})]$ with
$f_{x}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ $=-\sin\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),$ (29)
$f_{y}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ $=\cos\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),$ (30)
in a superconductor at (b) $V^{\mathrm{L}}_{0}=+1$, (c) $+2$, and (d) $+3$. In the presence of two mirror vortices, we obtain $J_{y}(y=-14.5a)=J(y=14.5a)=0$ with
$J_{y}(y)$ $=f_{y}(y)-f_{y}(y\pm a),$ (31)
as the boundary condition.
Figure 8: Amplitude of the first harmonics of the Josephson current versus the vortex core positions for different vortex winding: (a,b) $V^{\mathrm{L}}_{0}=1$, (c,d) $V^{\mathrm{L}}_{0}=2$, (e,f) $V^{\mathrm{L}}_{0}=3$. $I_{1}$ and $J_{1}$ indicate the odd- and even-parity first harmonics of the supercurrent evaluated by the direct Cooper pairs tunneling. Gray-dotted and dashed lines stand for the vortex core positions, $y^{\mathrm{L}}_{0}$ and $x^{\mathrm{L}}_{0}$, evaluated in Fig. 5. The aspect ratio is $\alpha=3/2$.
By taking into account the modification of the superconducting phase due to the image vortices, in Figs. 7 (b)-(m), we present the rectification amplitude for different values of the vortex winding and by varying the vortex core position assuming $\alpha=3/2$ as representative cases for the aspect ratio of the superconducting lead. We also analyze the rectification with the boundary condition [Figs. 7 (b)-(m)] as we worked in Sect. III. For a conventional vortex with winding $V^{\mathrm{L}}_{0}=1$ in Figs. 2 (b,c), even with the boundary condition, the rectification has a similar tendency without the boundary condition [Figs. 2 (b,c)]. Other vortex core positions away from the symmetry lines also give a negligible rectification. To maximize the supercurrent rectification, the amplitude of the first harmonics in the current phase relation also plays a main role [Figs. 7 (d)(e)], as we discussed in Sect. III.
For the even high-winding, $V^{\mathrm{L}}_{0}=2$, we find that a small amplitude of the rectification occurs [Figs. 7 (f,g)] because of the large value of $I_{1}$ compared to $I_{2}$ and $J_{1}$ [Figs. 7 (h,i)]. For $V^{\mathrm{L}}_{0}=3$, we also observe that the maximal values of the rectification amplitude $\sim 30\%$ occur for the vortex core position that is near (away from) the mirror lines of the superconducting lead [Figs. 7 (j) (Figs. 7 (k))]. Their profile is qualitatively similar to those for $V^{\mathrm{L}}_{0}=1$ [Figs. 7 (d,e,j,k)], due to the boundary condition.
As discussed in Sect. III, we can also derive the analytical expression of the Josephson current:
×$\times[\sin\varphi\cos\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})+\cos\varphi\sin\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})],$
by $\varphi^{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})\rightarrow\tilde{\varphi}^{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$. Likewise, the effective first harmonics components $\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0})$ and $\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$ of $\tilde{I}$ are given by
$\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\cos\tilde{\varphi}_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),$ (33)
$\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum^{0}_{j_{x},j_{y}}\sin\tilde{\varphi}_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}).$ (34)
We find that their nodal lines can cross in different sites depending on the even and odd winding number [Fig. 8]. Due to the boundary condition, the same odd functions are qualitatively obtained for the odd winding number $V^{\mathrm{L}}_{0}=1,3$ [Figs. 8 (a,b,e,f)]. Even functions without (with) the nodal structure for $\bar{I}_{1}$ ($\bar{J}_{1}$) are demonstrated at $V^{\mathrm{L}}_{0}=2$ [Figs. 8 (c,d)].
V Conclusions
We have demonstrated that nonreciprocal supercurrents can be exploited to distinguish physical configurations with vortices having winding number equal to one, coexistence of vortices and antivortices or the occurrence of a giant vortex. A suitable platform to observe these effects can be based on Josephson junctions made of nanoislands in the presence of an applied magnetic field [13]. It is known that giant vortices can be induced by magnetic field for superconductors with coherence length much smaller than the magnetic penetration depth (e.g. Pb or Nb based nanostructures) [13]. Then, starting from a low magnetic field configuration with a vortex having, for instance, winding $V^{\mathrm{L}}_{0}=1$ nearby the center of the superconducting lead, the transition to a higher winding vortex state by the increase of the magnetic field will be accompanied to a sizable variation of the rectification. Detection of such transitions can be used to probe the high-winding vortex phase. Along this line, one can also envision a field-free platform [53] with circular current flow inducing high-winding vortices that concomitantly yield supercurrent rectification. Moreover, since a generic phase vortex texture can be expanded in harmonics by employing vortex configurations with different windings, then, in principle our results provide a toolkit to design the supercurrent rectification for a wide variety of superconducting phase patterns.
VI Acknowledgments

### v5

@@ Introduction
Conclusions
We have demonstrated that nonreciprocal supercurrents are achieved in the presence of high-winding vortex and multiple vortices. The resulting behavior contains distinct features that can be exploited to distinguish physical configurations with vortices having winding number equal to one, coexistence of vortices and antivortices, or the occurrence of a giant vortex. In particular, we uncover the spatial profile of the supercurrent rectification with respect to the vortex core position. The rectification pattern depends on the number of windings, with an increase in the winding number producing more complex nodal structures. We find that for $V^{\mathrm{L}}_{\mathrm{0}}=1$, a single sign change line exists, while for 2 and 3, multiple lines and nodal points appear, demonstrating how winding number shapes the spatial rectification behavior. Our findings indicate that the achieved vortex diodes do not exhibit a high rectification efficiency. However, in this context, unlike semiconducting diodes, there are no fundamental reasons to exclude the use of superconducting diodes in superconducting electronics and quantum circuitry, even if their rectification efficiency does not reach $100\%$. Recent reports indeed highlighted the potential use of superconducting diodes in diverse applications [63, 64, 65], demonstrating that even with low rectification efficiencies, superconducting diodes can effectively perform functions such as alternating current (AC) to direct current (DC) conversion and rectification. In particular, one of the primary applications of superconducting diodes involves converting AC to DC at low temperatures to generate stable and adjustable DC bias currents from radiofrequency signals. They achieve this with a rectification efficiency below $50\%$ exploiting vortex dynamics, with the nonreciprocal critical current that arises from the asymmetric expulsion of vortices from the superconducting nanostructure. In our vortex diode design, instead, we assume that the nonreciprocal supercurrent can be controlled by the vortex position within the junction, without the need to move the vortex itself, where one can achieve rectification amplitudes of the order of $30\%$. In this framework, it is worth pointing out that, with respect to the design of vortex diodes and control knobs, vortices can be manipulated (e.g. displaced, introduced, or removed) by magnetic field [39, 66, 67], current [68, 69, 70], light [71, 72], and mechanical strain [73].
Figure 6: A schematic illustration of a physical setup designed to probe nonreciprocal phenomena with high-winding vortices. An external current source ($I_{1}$) injects current into a superconductor labeled $S$ (depicted in red). The circulating current influences a nearby proximitized superconductor, $S^{{}^{\prime}}$ (shown in orange), leading to the nucleation of an electrically controlled vortex within it [74]. A supercurrent, $I_{2}$, flows through a weak link based on $S^{{}^{\prime}}$ superconducting leads, where an electrically controlled vortex exists on one side of the junction. The nonreciprocal behavior of $I_{2}$ can be exploited to identify the nature of the induced vortex.
We would like to discuss how the achieved results depend on system parameters like temperature, junction interface, and disorder. Within our model description, the effects of temperature are to substantially reduce the amplitude of the superconducting order parameter. Hence, we do not expect qualitative changes but rather a modification of the amplitude of the supercurrent. Regarding the disorder, as shown in Section D of the Supplementary Information, we find that the rectification amplitude is not much altered by the local disorder potential. What is more relevant is the junction transparency as it enters to modify the second harmonic component of the supercurrent, as predicted by the Kulik-Omelyanchuk theory, for a superconducting weak link. We have examined the rectification amplitude for a representative vortex configuration in terms of the junction transparency for a representative vortex configuration (see Supplementary Information - Section D). Our results confirm that the rectification amplitude gets suppressed when the superconducting junction is brought into the tunneling regime.
It is also interesting to comment on the possibility of having a coexistence of vortex states and finite momentum pairing, a physical scenario that might lead to nontrivial effects for the superconducting vortex diode. However, it is unlikely that the vortex phase enables the formation of finite momentum pairing. This can happen for large values of the applied magnetic field when a vortex lattice can coexist with a finite momentum pairing of the FFLO type [75]). This is, however, in a regime of an applied magnetic field, which is beyond the examined cases in our paper because the coexistence occurs close to the upper critical field. Our study refers to a small applied magnetic field where only a few vortices nucleate into the superconductor.
We also would like to mention that our results refer to short superconducting junctions. In short Josephson junctions, the current phase relation’s structure and strong coupling can enhance diode efficiency by creating pronounced asymmetries. Instead, long Josephson junctions (LJJs), with their complex phase dynamics and fluxon motion, can manifest different mechanisms for the diode effect. The fluxonium diode, for instance, employs a control line to induce magnetic field asymmetry, though its performance remains not fully tested [76, 77]. Another concept involves a single annular junction with a control line requiring precise flux insertion [78]. Recent advances include asymmetric inline LJJs demonstrating sizable superconducting diode effects [79]. A suitable platform to observe these effects can be based on Josephson junctions made of nanoislands in the presence of an applied magnetic field [13]. It is known that giant vortices can be induced by a magnetic field for superconductors with coherence length much smaller than the magnetic penetration depth (e.g. Pb or Nb based nanostructures) [13]. In particular, as demonstrated by the theoretical and experimental results in Ref.  [13], for systems with a size approximately five times the coherence length, the application of a magnetic field on the order of tens of millitesla can induce transitions involving changes in the vortex winding configurations. Our calculations are based on a system size and coherence length that align with this analysis. Then, starting from a low magnetic field configuration with a vortex having, for instance, winding $V^{\mathrm{L}}_{0}=1$ nearby the center of the superconducting lead, the transition to a higher winding vortex state by the increase of the magnetic field will be accompanied by a sizable variation of the rectification. Detection of such transitions can be used to probe the high-winding vortex phase.
Along this line, one can also envision a field-free platform [74] with circular current flow inducing high-winding vortices that concomitantly yield supercurrent rectification (see Fig. 6). The described setup provides a potential approach to investigating nonreciprocal phenomena through the manipulation of high-winding vortices. By utilizing an external current to induce a vortex in a proximitized superconductor and examining the resulting supercurrent behavior across a weak link, this configuration offers a means to characterize and understand the underlying physics of nonreciprocity in superconducting systems with nontrivial vortex states. Such insights could pave the way for advanced superconducting devices with directional control and enhanced functionalities. Moreover, since a generic phase vortex texture can be expanded in harmonics by employing vortex configurations with different windings, then, in principle, our results provide a toolkit to design the supercurrent rectification for a wide variety of superconducting phase patterns.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Acknowledgments